OpenStudy (anonymous):

can you add radicals like 74^1/2 + 13^1/2

OpenStudy (anonymous):

\[\sqrt{74} + \sqrt{13}\] thats what it looks like

OpenStudy (anonymous):

no

OpenStudy (anonymous):

they are not like terms, and \(\sqrt{74}\) is in simplest radical form, so there is nothing you can do with it

OpenStudy (anonymous):

ok cool